OpenStudy (anonymous):

What is the largest equilateral triangle that can be inscribed in a regular pentagon with side length 1?

OpenStudy (eyust707):

my intuition says it will look something like this: 

OpenStudy (eyust707):

theres probably an easier way but the only thing i can think of is to create a system of equations that that must satisfy

OpenStudy (anonymous):

could you walk me through it?

OpenStudy (eyust707):

i honestly dont know how exactly... but i could exaplin to you my thought process

OpenStudy (eyust707):

what level math is this for?

OpenStudy (anonymous):

haha really high level

OpenStudy (eyust707):

yea its not an easy problem

OpenStudy (eyust707):

lets start with the angles

OpenStudy (eyust707):

each one is 72 degrees

OpenStudy (anonymous):

http://www.twiddla.com/806032

OpenStudy (anonymous):

This is no where near high level mathematics! Hint: http://www.drking.org.uk/hexagons/misc/deriv1.html (This is for hexagon)

OpenStudy (eyust707):

that doesnt actually prove thats the biggest one but if we assume so its just simple trig

OpenStudy (anonymous):

how did you know that the other angle is 72?

OpenStudy (eyust707):

360/5

OpenStudy (dumbcow):

no it would be 540/5 sum interior = 180(n-2)

OpenStudy (eyust707):

ahh yes 72 is the central angles

OpenStudy (eyust707):

clearly its bed time for me

OpenStudy (dumbcow):

 now use law of sines \[x = \frac{\sin 108}{\sin 48}\]